OpenStudy (anonymous):

what do you do to evaluate sec -60 degrees

OpenStudy (anonymous):

sec(-60)=sec60 =2 as sec(-x)=secx

OpenStudy (anonymous):

so the answer is 2 right

jimthompson5910 (jim_thompson5910):

you can add to it sec(60) = 1/cos(60) sec(60) = 1/(1/2) sec(60) = (1/1)/(1/2) sec(60) = (1/1)*(2/1) sec(60) = 2 So that's how matricked got the answer of 2

OpenStudy (anonymous):

gotcha thanks!

jimthompson5910 (jim_thompson5910):

np